A constant which has no unit
1
Expert's answer
2020-05-12T09:59:01-0400

Equilibrium constant K.


Suppose, there is a chemical reaction A + B = C + D


K = ([C][D]) / ([A][B]), where [A], [B], [C] and [D] - equilibrium concentrations of the substances A, B, C and D, measured in mol/L. As everithing is measured in mol/L, then K = (mol2/L2) / (mol2/l2), and so it has no unit, only numerical value.